OpenStudy (anonymous):

dv= 8x^2dx, what is v=dx? Please help

OpenStudy (anonymous):

Don't really understand this, let's integrate so v = 8x^3/3 + C Does that help?

OpenStudy (anonymous):

if i may ask how would you know..i dont know how i would get 8x^2dx

OpenStudy (anonymous):

"i dont know how i would get 8x^2dx" ???? You gave this in the question.....

OpenStudy (anonymous):

yes but im asking how would you derivrate 8/3 x^3 to 8x^2?

OpenStudy (anonymous):

Ah, OK..... v = 8x^3/3 + C so dv/dx = 3* 8x^(3-1) / 3

OpenStudy (anonymous):

In general, ax^n dx = n*a * x^(n-1)

OpenStudy (anonymous):

oh ok i see thanks so much!!

OpenStudy (anonymous):

ur welcome